# Black Hole Superradiance of Interacting Multi-Field

### v1

@@ ACKNOWLEDGMENTS
J. Z. is supported by the National Natural Science Foundation of China (NSFC) under Grants No. E414660101 and No. 12147103, and the Fundamental Research Funds for the Central Universities under Grants No. E4EQ6604X2 and No. E3ER6601A2. Y.-S. P. is supported by National Key Research and Development Program of China under Grant No.2021YFC2203004 and the NSFC under Grant No.12475064.
Appendix A Stability of quasi-equilibrium
@@
$\dot{\epsilon}$ $=\gamma_{sr}(a)\epsilon-\gamma_{k}\epsilon^{k+1}$ (59)
$\dot{a}$ $=-\gamma_{sr}(a)\epsilon\,$,
where $k\geq 1,k\in\mathbb{Z}$ and $\gamma_{k}$ is a constant. Assuming the system is in equilibrium, we can substitute the equilibrium occupation number $\epsilon_{\rm eq}(a)=(\gamma_{sr}/\gamma_{k})^{1/k}$ into Eq. (59) and get
$\mathrm{d}\eta=-\frac{\mathrm{d}a}{\epsilon_{\rm eq}^{k+1}}\,$, (60)
where $\eta=\gamma_{k}\tilde{t}$ is the rescaled time. Integrating both sides of the equation and expand $a$ around $a_{\rm crit}$, we have
$\eta=k(a-a_{\rm crit})^{-\tfrac{1}{k}}\left(\frac{\alpha^{8}}{24\gamma_{k}}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)^{-\tfrac{1+k}{k}}\,$, (61)
@@
$\epsilon_{\rm eq}=\left(\frac{a-a_{\rm crit}}{\gamma_{k}}\frac{\alpha^{8}}{24}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)^{\tfrac{1}{k}}\,.$ (62)
@@
$a-a_{\rm crit}$ $=\left(\frac{k}{\eta}\right)^{k}\left(\frac{24\gamma_{k}}{\alpha^{8}}\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\right)^{k+1}$ (63)
@@
which means a typical timescale for restoring equilibrium is $\tau_{re}\sim(k\epsilon_{\rm eq}^{k})^{-1}$. On the other hand, $\epsilon_{\rm eq}$ varies on a time scale of
$\tau_{var}\sim\left|\frac{\epsilon_{\rm eq}}{\epsilon^{\prime}_{\rm eq}}\right|=\left|\frac{\epsilon_{\rm eq}}{\frac{{\rm d}\epsilon_{\rm eq}}{{\rm d}a}a^{\prime}}\right|=\left|\frac{\mathrm{d}\epsilon_{\rm eq}}{\mathrm{d}a}\epsilon_{\rm eq}^{k}\right|^{-1}\,.$ (65)
Therefore, we expect that equilibrium breaks down when $\tau_{re}\gtrsim\tau_{var}$, as the evolution of real occupation number $\epsilon$ can not catch up with the change of $\epsilon_{\rm eq}$. In case A, we have $k=1$, and $\tau_{re}<\tau_{var}$ is always satisfied when $a$ is close to $a_{\rm crit}$. For $k\geq 1$, there could be threshold in terms of black hole spin $a_{th}$, below which one has $\tau_{re}>\tau_{var}$, indicating the equilibrium is unstable. After the black hole spin drops below $a_{th}$, the power law decay of $\epsilon$ transforms from $t^{-1}$ to $t^{-1/k}$.

### v2

@@ ACKNOWLEDGMENTS
J. Z. is supported by the National Natural Science Foundation of China (NSFC) under Grants No. E414660101 and No. 12347103, and the Fundamental Research Funds for the Central Universities under Grants No. E4EQ6604X2 and No. E3ER6601A2. Y.-S. P. is supported by National Key Research and Development Program of China under Grant No.2021YFC2203004 and the NSFC under Grant No.12475064.
Appendix A Ultralight scalar fields in the weak gravity limit of the Kerr black hole background
In this appendix, we shall discuss the solution to the Klein-Gordon equation in the weak gravity limit of the Kerr black hole background. Let us start with the Klein-Gordon equation in the Boyer-Lindquist coordinates
$\left(-\Box+\mu^{2}\right)\psi=0\,.$ (64)
We only consider the case that $\alpha\equiv GM\mu\ll 1$, where $M$ is the black hole mass. Far from the black hole horizon, the D’Alembert operator $\Box$ can be expand in powers of $GM/r$. To the first order of $GM/r$, Eq. (64) can be written in the form of
$\left[\partial_{t}^{2}-\nabla^{2}+\mu^{2}-\frac{2GM}{r}\left(\mu^{2}+\hat{K}^{2}\right)\right]\psi\simeq 0\,$, (65)
where $\nabla^{2}=\partial_{r}^{2}+\tfrac{2}{r}\partial_{r}-\tfrac{\hat{L}^{2}}{r^{2}}$ is the Laplace operator in Minkowski space and
$\hat{K}^{2}$ $\equiv-\frac{1}{r}\partial_{r}-\partial_{r}^{2}-\mu^{2}-\partial_{t}^{2}$ (66)
$=-\frac{3}{r}\partial_{r}+\frac{\hat{L}^{2}}{r^{2}}-2\partial_{r}^{2}$
with $\hat{L}^{2}$ being the total angular momentum operator. In the second line of above equation, we have used the equation of free $\psi$ at zeroth order of $GM/r$ to eliminate the $\partial_{t}^{2}$ operator. For non-relativistic states, the wavenumber $k$ is much less than $\mu$. Therefore, $\hat{K}^{2}\sim k^{2}$ can be neglected, and we have
$\left[\partial_{t}^{2}-\nabla^{2}+\mu^{2}-\frac{2GM\mu^{2}}{r}\right]\psi\simeq 0\,$, (67)
for non-relativistic states in the weak gravity limit.
We shall proceed in the frequency domain by considering
$\psi=e^{-i\omega t}u(\mathbf{r})\,$, (68)
and hence
$\left(-\omega^{2}+\mu^{2}-\nabla^{2}-\frac{2\alpha\mu}{r}\right)u(\mathbf{r})\simeq 0\,.$ (69)
$\left(-\nabla^{2}-\frac{2\alpha\mu}{r}\right)v(\mathbf{r})\simeq(\sigma^{2}-\nu^{2})v(r)\,.$ (70)
The general solution to Eq. (70) are well known in the literature. By imposing the regular boundary condition, one can select a set of relevant eigenstates $\{u_{n\ell m}\}$ and $\{u_{k\ell m}\}$, which satisfy
$\left(-\nabla^{2}-\frac{2\alpha\mu}{r}\right)u_{n\ell m}$ $=-k_{n}^{2}u_{n\ell m},\quad\frac{k_{n}}{\alpha\mu}=\frac{1}{n},n\in\mathbb{N}$ (71)
$\left(-\nabla^{2}-\frac{2\alpha\mu}{r}\right)u_{k\ell m}$ $=k^{2}u_{k\ell m},\quad\frac{k}{\alpha\mu}\in(0,\infty)$
respectively. Specifically, the eigenfunctions are given by $u_{n(k)\ell m}=R_{n(k)\ell m}(r)Y_{\ell m}(\theta,\phi)$ with
$R_{n\ell}=$ $\sqrt{\left(\frac{2}{nr_{B}}\right)^{3}\frac{(n-\ell-1)!}{2n(n+\ell)!}}e^{-\tfrac{r}{nr_{B}}}\left(\frac{2r}{nr_{B}}\right)^{\ell}L^{2\ell+1}_{n-\ell-1}\left(\frac{2r}{nr_{B}}\right),$ (72)
$R_{k\ell}=$ $\frac{2ke^{\pi/(2kr_{B})}|\Gamma(\ell+1-i/(kr_{B}))|}{(2\ell+1)!}(2kr)^{\ell}$
$\times e^{-ikr}M\left(i/(kr_{B})+\ell+1,2\ell+2,2ikr\right)$
where $r_{B}\equiv 1/\alpha\mu$, $Y_{\ell m}$ is the spherical harmonics, $L^{2\ell+1}_{n-\ell-1}$ is the generalized Laguerre polynomial of degree $n-\ell-1$ and $M(i/(kr_{B})+\ell+1,2\ell+2,2ikr)$ is Kummer’s function of the first kind. $R_{k\ell}$ are also known as the Coulomb wave function $F_{\ell}(\eta,r)$. Given the asymptotical behavior at spacial infinity, one can find that $\{u_{n\ell m}\}$ correspond to bound states while $\{u_{k\ell m}\}$ correspond to unbound states. The eigenfunctions obey the following orthogonality condition,
$\int u_{n_{1}\ell_{1}m_{1}}u_{n_{2}\ell_{2}m_{2}}^{*}\mathrm{d}^{3}\mathbf{r}=\delta_{n_{1}n_{2}}\delta_{\ell_{1}\ell_{2}}\delta_{m_{1}m_{2}}$ (73)
$\int u_{k_{1}\ell_{1}m_{1}}u_{k_{2}\ell_{2}m_{2}}^{*}\mathrm{d}^{3}\mathbf{r}=2\pi\delta(k_{1}-k_{2})\delta_{\ell_{1}\ell_{2}}\delta_{m_{1}m_{2}}$
$\int u_{n\ell_{1}m_{1}}u_{k\ell_{2}m_{2}}^{*}\mathrm{d}^{3}\mathbf{r}=0\,.$
In the literature, the bound states are also often described with $\bar{\psi}$} defined by the non-relativistic ansatz (6). By substituting ansatz (6) into Eq. (67) and keeping leading order terms in $\alpha$, we can find that $\bar{\psi}$} satisfies
$\left(i\partial_{t}+\frac{\nabla^{2}}{2\mu}+\frac{\alpha}{r}\right){\bar{\psi}}\simeq 0\,$, (74)
and the eigenstates are given by
${\bar{\psi}}_{n\ell m}=e^{-i\xi_{n\ell m}t}u_{n\ell m}(\mathbf{r})$ (75)
with $\xi_{n\ell m}\simeq\omega_{n\ell m}-\mu$. For relativistic states, one should take $\hat{K}^{2}$ into account. Nevertheless, when we calculation the radiation power of the $3{\bar{\psi}}$-process, we take the flat-space approximation for simple, in which case the eigenfunction of the unbound states can be obtained by taking $u_{k\ell m}$ to the limit of $kr_{B}\rightarrow\infty$, and the radial function $R_{k\ell}$ reduces to spherical Bessel function.
Appendix B Green function method
In this appendix, we shall show how to solve the inhomogeneous Klein-Gordon equation [^5] Here we use the $\psi$ field to demonstrate the method, while in the main text, the Green function method is applied to solve Eq. (17) and Eq. (34) for the $\varphi$ field.
$\left(\Box-\mu^{2}\right)\psi(x)=s(x)\,.$ (76)
on the Kerr background with the Green function method. We shall set the Newton’s constant $G=1$ in this appendix to avoid confusion with the Green function. In the Boyer-Lindquist coordinates, the equation is separable with
$\psi(x)=\int d\omega e^{-i\omega t}R_{\omega\ell m}(r)S_{\omega\ell m}(\theta)e^{im\phi}\,$ (77)
where $S_{\omega\ell m}(\theta)$ is the spheroidal harmonic function. For simplicity, we shall consider
$s(x)=F(r)\,S_{\Omega\ell m}(\theta)\,e^{-i\Omega t}+c.c.\,$, (78)
such that we can focus on the radial function which stratifies
$\frac{\mathrm{d}}{\mathrm{d}r}\left(\Delta\frac{\mathrm{d}R}{\mathrm{d}r}\right)+\Bigg[$ $\frac{(r^{2}+a^{2})^{2}\omega^{2}-4Mam\omega r+m^{2}a^{2}}{\Delta}$
$-\left(a^{2}\omega^{2}+\mu^{2}r^{2}+\Lambda_{\ell m}\right)\left.\Bigg]\right|_{\omega=\Omega}R=I(r)\,$, (79)
where $\Delta=r^{2}-2Mr+a^{2}$ and $\Lambda_{\ell m}$ is the eigenvalue of angular Teukolsky equation, and we have suppressed the subscription of the radial function for short.
B.1 The radial Green function
The radial function can be solved with the radial Green function,
$R(r_{*})=\int_{-\infty}^{+\infty}I(r^{\prime}_{*})\,G_{\ell m\omega=\Omega}(r_{*},r^{\prime}_{*})\,\mathrm{d}r^{\prime}_{*}$ (80)
where $r_{*}$ is the tortoise coordinate. The Green function can be construct as
$G_{\ell m\omega}(r_{*},r^{\prime}_{*})=\frac{f_{H}(r_{*<};\omega)f_{+}(r_{*>};\omega)}{\mathcal{W}(f_{H},f_{+})}$ (81)
where $r_{*<}\equiv\min\{r^{\prime}_{*},r_{*}\}$ $r_{*>}\equiv\max\{r^{\prime}_{*},r_{*}\}$ and $\mathcal{W}(f_{H},f_{+})\equiv f_{H}\partial_{r_{*}}f_{+}-f_{+}\partial_{r_{*}}f_{H}$ is the Wronskian of the two independent homogeneous solutions $f_{H}$ and $f_{+}$. The homogeneous solutions are defined by their asymptotic properties
[…]
where $k_{H}=\omega-m\Omega_{H}$, $k_{\infty}=\sqrt{\omega^{2}-\mu^{2}}$. Another useful solution is
$f_{-}\sim\begin{cases}C_{\mathrm{in}}(\omega)e^{-ik_{H}r_{*}}+C_{\mathrm{out}}(\omega)e^{ik_{H}r_{*}}\quad&r_{*}\rightarrow-\infty\\ e^{-ik_{\infty}r_{*}}\quad&r_{*}\rightarrow+\infty\end{cases},$ (83)
which describes the in-going wave at infinity. Since there are only two independent solutions, the above solutions are related by
$f_{H}=A_{\mathrm{in}}(\omega)f_{-}+A_{\mathrm{out}}(\omega)f_{+}\,$, (84)
and the corresponding Wronskians have
$\mathcal{W}(f_{H},f_{+})=2ik_{\infty}A_{\mathrm{in}}(\omega)\,$, (85)
$\mathcal{W}(f_{H},f_{-})=-2ik_{\infty}A_{\mathrm{out}}(\omega)\,.$ (86)
B.2 Relation to the eigenstates
Figure 13: Analyticity of $\zeta(z)$ and the integral contour in integral (87). The branch cut is chosen at $(-\infty,-\mu]$ and $\mu,\infty)$. The left plot shows the principal Riemann sheet, and the right plot shows the second Riemann sheet. There are poles at $\pm\Omega$ (the red dots), and at the zeros of Wronskian (the black dots). The bound state poles (${\rm Re}\omega<\mu$) are on the principle Riemann sheet, while the qausi-normal mode poles are on the second Riemann sheet.
Now we show that Eq. (80) can be expanded with the eigenfunctions of homogeneous radial equation, i.e. Eq. (B) with $I=0$. We shall show this by constructing an integral
$\frac{1}{2\pi i}\oint_{C}\zeta(z)\,\mathrm{d}z\,$ (87)
where
$\zeta(z)=\frac{2z}{z^{2}-\Omega^{2}}\int_{-\infty}^{+\infty}I(r^{\prime}_{*})G_{\ell mz}(r_{*},r^{\prime}_{*})\mathrm{d}r^{\prime}_{*}\,.$ (88)
Recall that $\Omega$ is the forced oscillation frequency of the source term $s(x)$. The integration contour and the analyticity of $\zeta(z)$ are shown in Fig. 13. Due to the multi-value of $k=\pm\sqrt{z^{2}-\mu^{2}}$, $\zeta(z)$ has two branch point $z=\pm\mu$ and we define the integral in the Riemann sheet of ${\rm Im}\,k>0$ to fit the out-going boudnary condition, which means the branch cut is $(-\infty,-\mu]$ and $\mu,\infty)$.
We will see in the following that the integral connects three parts: the residues at $z=\pm\Omega$ which correspond to Eq. (80), the residues at $\omega_{n}$ which correspond to the contribution from bound states, and the branch cut which corresponds to the contribution from scattering states.
According to the residue theorem, we have
$\frac{1}{2\pi i}\oint_{C}\zeta(z)\mathrm{d}z=\sum_{z_{0}=\pm\Omega}{\rm Res}\left(\zeta,z_{0}\right)+\sum_{n}{\rm Res}\left(\zeta,\omega_{n}\right)$ (89)
where $\omega_{n}$ are the relevant poles from $G_{\ell mz}(r_{*},r^{\prime}_{*})$. It is straightforward that the residues at $\pm\Omega$ give $R(r_{*})$ obtained from Eq. (80). Besides $\pm\Omega$, the zeros of Wronskian also lead to other poles. At these poles, we have $f_{H}=A_{\rm out}f_{+}\propto f_{n}$, which are the eigenfunctions of bound states and quasi-normal modes by definition. Specifically, the residue can be written
${\rm Res}\left(\zeta,\omega_{n}\right)=$ $\frac{\int_{-\infty}^{+\infty}I(r^{\prime}_{*})f_{n}(r^{\prime}_{*})\mathrm{d}r^{\prime}_{*}}{\omega_{n}^{2}-\Omega^{2}}f_{n}(r_{*})\,$, (90)
where $f_{n}$ is normalized by
$f_{n}(r_{*})f_{n}(r_{*})\equiv$ $2\omega_{n}{\rm Res}\left(G_{z}(r_{*},r^{\prime}_{*}),\omega_{n}\right)$ (91)
$=$ $-\frac{2\omega_{n}f_{H}(r_{*};\omega_{n})f_{H}(r_{*};\omega_{n})}{2ik_{\infty}A_{\mathrm{out}}\partial_{\omega_{n}}A_{\mathrm{in}}(\omega_{n})}\,$,
such that it approaches to the normalized eigenfunctions discussed in App. A in the weak gravity limit.
On the other hand, since $G_{\ell mz}(r_{*},r^{\prime}_{*})\sim 1/\sqrt{z^{2}-\mu^{2}}$ when $|z|\rightarrow\infty$, the integral going around the circle vanishes, leaving only the contribution from the branch cut. Around the right branch cut, we have
$I_{\mathrm{cut}}=$ $\int_{\mu}^{\infty}\zeta(z+i\epsilon)-\zeta(z-i\epsilon)\mathrm{d}z$ (92)
$=$ $\int_{\mu}^{\infty}\mathrm{d}z\frac{2z}{z^{2}-\Omega^{2}}\int_{-\infty}^{+\infty}\mathrm{d}r^{\prime}_{*}I(r^{\prime}_{*})f_{H}(r_{*<};z)$
$\times\left(\frac{f_{+}(r_{*>};z+i\epsilon)}{\mathcal{W}(z+i\epsilon)}-\frac{f_{+}(r_{*>};z-i\epsilon)}{\mathcal{W}(z-i\epsilon)}\right)$
$=$ $\int_{\mu}^{\infty}\mathrm{d}z\frac{2z}{z^{2}-\Omega^{2}}\int_{-\infty}^{+\infty}\mathrm{d}r^{\prime}_{*}I(r^{\prime}_{*})f_{H}(r_{*<};z)$
$\times\left(\frac{f_{+}(r_{*>};z)}{\mathcal{W}[f_{H}(z),f_{+}(z)]}-\frac{f_{-}(r_{*>};z)}{\mathcal{W}[f_{H}(z),f_{-}(z)]}\right)$
$=$ $\int_{\mu}^{\infty}\mathrm{d}z\frac{2z}{z^{2}-\Omega^{2}}\int_{-\infty}^{+\infty}\mathrm{d}r^{\prime}_{*}I(r^{\prime}_{*})f_{H}(r_{*<};z)\frac{A_{\mathrm{out}}f_{+}+A_{\mathrm{in}}f_{-}}{2ik_{\infty}A_{\mathrm{in}}A_{\mathrm{out}}}$
$=$ $\int_{0}^{\infty}\mathrm{d}k\frac{1}{z^{2}(k)-\Omega^{2}}\int_{-\infty}^{+\infty}\mathrm{d}r^{\prime}_{*}I(r^{\prime}_{*})\frac{f_{H}(r_{*<};z)f_{H}(r_{*>};z)}{iA_{\mathrm{in}}A_{\mathrm{out}}}.$
Eventually, by equating the above result with the result from the residue theorem, we have
$R(r_{*})=$ $\sum_{n}\frac{\int_{-\infty}^{+\infty}I(r^{\prime}_{*})f_{n}(r^{\prime}_{*})\mathrm{d}r^{\prime}_{*}}{\omega_{n}^{2}-\Omega^{2}}f_{n}(r_{*})$ (93)
$+\int_{0}^{\infty}\frac{\mathrm{d}k}{2\pi}\frac{\int_{-\infty}^{+\infty}\mathrm{d}r^{\prime}_{*}I(r^{\prime}_{*})f_{H}(r^{\prime}_{*};z)}{z^{2}(k)-\Omega^{2}}\frac{f_{H}(r_{*};z)}{A_{\mathrm{in}}A_{\mathrm{out}}}\,.$
For $s(x)=\tilde{s}(\mathbf{r})e^{-i\Omega t}+c.c.$ with general angular dependence, one can consider the full Green function [68]
$G_{\rm ret}(x,x^{\prime})=$ $\frac{2}{\sqrt{r^{2}+a^{2}}\sqrt{r^{\prime 2}+a^{2}}}\int d\omega e^{-i\omega(t-t^{\prime})}$
$\times\sum_{m}e^{im(\phi-\phi^{\prime})}\sum_{l}S_{\omega\ell m}(\theta)S_{\omega\ell m}^{*}(\theta^{\prime})$
$\times G_{\ell m\omega}(r,r^{\prime})$ (94)
and use Eq. (93) when computing the radial part. Specifically, in the weak gravity limit, we have $A_{\mathrm{in}}A_{\mathrm{out}}=1$, $f_{n}(r)S_{\omega\ell m}(\theta)e^{im\phi}$ and $f_{H}(r)S_{\omega\ell m}(\theta)e^{im\phi}$ reduce to $u_{n\ell m}(\mathbf{r})$ and $u_{k\ell m}(\mathbf{r})$ discussed in App. A respectively. Therefore, the solution to Eq. (19) can be given by Eq. (LABEL:eq:expPhi).
Appendix C Stability of quasi-equilibrium
@@
$\dot{\epsilon}$ $=\gamma_{sr}({\tilde{a}})\epsilon-\gamma_{k}\epsilon^{k+1}$ (95)
$\dot{{\tilde{a}}}$ $=-\gamma_{sr}({\tilde{a}})\epsilon\,$,
where $k\geq 1,k\in\mathbb{Z}$ and $\gamma_{k}$ is a constant. Assuming the system is in equilibrium, we can substitute the equilibrium occupation number $\epsilon_{\rm eq}({\tilde{a}})=(\gamma_{sr}/\gamma_{k})^{1/k}$ into Eq. (95) and get
$\mathrm{d}\eta=-\frac{\mathrm{d}{\tilde{a}}}{\epsilon_{\rm eq}^{k+1}}\,$, (96)
where $\eta=\gamma_{k}\tilde{t}$ is the rescaled time. Integrating both sides of the equation and expand $\tilde{a}$} around ${\tilde{a}}_{\rm crit}$, we have
$\eta=k({\tilde{a}}-{\tilde{a}}_{\rm crit})^{-\tfrac{1}{k}}\left(\frac{\alpha^{8}}{24\gamma_{k}}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)^{-\tfrac{1+k}{k}}\,$, (97)
@@
$\epsilon_{\rm eq}=\left(\frac{{\tilde{a}}-{\tilde{a}}_{\rm crit}}{\gamma_{k}}\frac{\alpha^{8}}{24}\frac{1+4\alpha^{2}}{1-4\alpha^{2}}\right)^{\tfrac{1}{k}}\,.$ (98)
@@
${\tilde{a}}-{\tilde{a}}_{\rm crit}$ $=\left(\frac{k}{\eta}\right)^{k}\left(\frac{24\gamma_{k}}{\alpha^{8}}\frac{1-4\alpha^{2}}{1+4\alpha^{2}}\right)^{k+1}$ (99)
@@
which means a typical timescale for restoring equilibrium is $\tau_{re}\sim(k\epsilon_{\rm eq})^{-1}$. On the other hand, $\epsilon_{\rm eq}$ varies on a time scale of
$\tau_{var}\sim\left|\frac{\epsilon_{\rm eq}}{\epsilon^{\prime}_{\rm eq}}\right|=\left|\frac{\epsilon_{\rm eq}}{\frac{{\rm d}\epsilon_{\rm eq}}{{\rm d}{\tilde{a}}}{\tilde{a}}^{\prime}}\right|=\left|\frac{\mathrm{d}\epsilon_{\rm eq}}{\mathrm{d}{\tilde{a}}}\epsilon_{\rm eq}^{k}\right|^{-1}\,.$ (101)
Therefore, we expect that equilibrium breaks down when $\tau_{re}\gtrsim\tau_{var}$, as the evolution of real occupation number $\epsilon$ can not catch up with the change of $\epsilon_{\rm eq}$. In case A, we have $k=1$, and $\tau_{re}<\tau_{var}$ is always satisfied when $a$ is close to ${\tilde{a}}_{\rm crit}$. For $k\geq 1$, there could be threshold in terms of black hole spin ${\tilde{a}}_{th}$, below which one has $\tau_{re}>\tau_{var}$, indicating the equilibrium is unstable. After the black hole spin drops below ${\tilde{a}}_{th}$, the power law decay of $\epsilon_{\rm eq}$ transforms from $t^{-1}$ to $t^{-1/k}$.